An employee earns exist25 per hour and 2 times that rate for all hours in excess of 40 hours per week. Assume that the employee worked 48 hours during the week. Assume further that the social security tax rate was 6.0%, the Medicare tax rate was 1.5%, and federal income tax to be withheld was exist 239.15.
a. Determine the gross pay for the week.
b Determine the net pay for the week.
Answer #1

SOLUTION

(A) Calculation of Gross pay

Average rate = $25 per hour

Overtime rate = $25*2 = $50 per hour

Gross pay = (40*$25) + (8 * $50) = $1,400

(B) Calculation of Net Pay

Social Security tax = $1,400 * 6% = $84

Medicare Tax = $1,400 * 1.5% = $21

Federal Income tax = $239.15

Net Pay = $1,400 - $84 - $21 - $239.15 = $1,055.85